OpenStudy (anonymous):

i need help on one last thing. it is... Determine the area required for a new landfill site with a projected life of 30 years for a population of 250,000 generating 2.02 kg per person per day. The density of the compacted waste is 470 kg /m3. The height of the landfill is not to exceed 15 m. Assume the land fill will be a rectangular cube.

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well find the amount of waste over the 30 year period...for the population divide it by the compaction factor... this will give the volume needed for the waste divide the volume by 15.... to get the area needed...

OpenStudy (anonymous):

don't know what i'md oing sorry

OpenStudy (campbell_st):

ok... so how much waste does the population generate in a day...?

OpenStudy (anonymous):

250000*30=7500000/470

OpenStudy (campbell_st):

250 000 * 30 = the population over 30 years how much waste do they create per day 250 000 * 2.02 then how much per year 250 000 *2.02 *365 how much in 30 years 250000 *2.02*365*30 = total waste now divide by the compaction factor to find the volume

OpenStudy (anonymous):

i'm getting11765425.53?

OpenStudy (campbell_st):

is that the volume...? if it is... divide by 15 to get the area.

OpenStudy (campbell_st):

I divided by 15 as that is the maximum height ... so you are actually finding the minimum area...

OpenStudy (anonymous):

if i divide by 15 i get 368650000

OpenStudy (campbell_st):

well the volume is correct... but the area seems incorrect... V = Area x height so Area = Volume/height \[A = \frac{11765425.53}{15}\]

OpenStudy (anonymous):

784361.702

OpenStudy (campbell_st):

thats what I got... so we may have both made the same mistakes... or its correct...

OpenStudy (campbell_st):

don't forget the area units in your answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (campbell_st):

glad to help